A sip of mass 3 into 10 power 7 kg is addressed it is 10.23 by applying a force of 5 into 10 ki power for and description of water is negligible then find speed of the ship

Grade:9

1 Answers

Arun
6 years ago
acceleration of ship a=5*104 N/ 3*107 KG
=5/3 *10-3 m/sec2
speed of the ship after movieng 3 m  
v2 =u2 +2aS
=0 +2* 5/3 *10-3   * (3)
= 10-2
v=.1 m/sec

Regards
Arun (askIITians forum expert)